# Universal constant order fluctuations for the cokernels of block lower triangular matrices

### v1

@@ A.1 The proof of (1)
$\mathbf{C}^{\prime}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\ddots&&&\\ \mathbf{I}&\mathbf{A}_{2}&\mathbf{0}&\ddots&\\ \mathbf{0}&\mathbf{I}&\mathbf{A}_{3}&\mathbf{0}&\ddots&\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ &\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-2}&\mathbf{0}&\ddots\\ &&\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-1}A_{k}&\mathbf{0}\\ &&\ddots&\ddots&\mathbf{0}&\mathbf{0}&\mathbf{I}\end{pmatrix}.$
@@
$\operatorname{cok}(\mathbf{C}^{\prime})\cong\operatorname{cok}\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\ddots&&&\\ \mathbf{I}&\mathbf{A}_{2}&\mathbf{0}&\ddots&\\ \mathbf{0}&\mathbf{I}&\mathbf{A}_{3}&\mathbf{0}&\ddots&\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ &\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-2}&\mathbf{0}&\\ &&\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-1}A_{k}\end{pmatrix}.$
@@ A.2 The proof of (3)
Using Lemmas 11 and 12, we have
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n},\,\langle\mathbf{g}\rangle=G,\\ \textrm{ $\mathbf{g}$ is not a code of distance $\delta n$}\end{subarray}}P(\mathbf{g})$ $\leq\sum_{\begin{subarray}{c}D>1\\ D\mid\#G\end{subarray}}\quad\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n},\,\langle\mathbf{g}\rangle=G,\\ \mathbf{g}\textrm{ has $\delta$-depth $D$}\end{subarray}}P(\mathbf{g})$
@@
Also, from Lemma 11, we can choose $\delta$ small enough so that we have
@@
Then given a choice of $\delta$ that satisfies the requirements above, using Lemma 10 we have a $c$ such that
@@
$\leq\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ \mathbf{g}\textrm{ is a code of dist. $\delta n$}\end{subarray}}\left|P(\mathbf{g})-|G|^{-n}\right|+\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n},\,\langle\mathbf{g}\rangle=G,\\ \textrm{ $\mathbf{g}$ is not a code of dist. $\delta n$}\end{subarray}}P(\mathbf{g})+\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ \mathbf{g}\textrm{ is not a code of dist. $\delta n$}\end{subarray}}|G|^{-n}$

### v2

@@ A.1 The proof of (2)
$\mathbf{C}^{\prime}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\ddots&&&\\ \mathbf{I}&\mathbf{A}_{2}&\mathbf{0}&\ddots&\\ \mathbf{0}&\mathbf{I}&\mathbf{A}_{3}&\mathbf{0}&\ddots&\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ &\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-2}&\mathbf{0}&\ddots\\ &&\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-1}\mathbf{A}_{k}&\mathbf{0}\\ &&\ddots&\ddots&\mathbf{0}&\mathbf{0}&\mathbf{I}\end{pmatrix}$.
@@
$\operatorname{cok}(\mathbf{C}^{\prime})\cong\operatorname{cok}\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\ddots&&&\\ \mathbf{I}&\mathbf{A}_{2}&\mathbf{0}&\ddots&\\ \mathbf{0}&\mathbf{I}&\mathbf{A}_{3}&\mathbf{0}&\ddots&\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ &\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-2}&\mathbf{0}&\\ &&\ddots&\ddots&\mathbf{I}&\mathbf{A}_{k-1}\mathbf{A}_{k}\end{pmatrix}$.
@@ A.2 The proof of (4)
Using Lemmas A.2 and A.3, we have
[…]
@@
Also, from Lemma A.2, we can choose $\delta$ small enough so that we have
@@
Then given a choice of $\delta$ that satisfies the requirements above, using Lemma A.1 we have a $c$ such that
@@
[…]